# The numeraire e-variable and reverse information projection

### v3

@@ 7 Composite alternatives and universal inference
To compare the universal inference e-variable with the numeraire $X^{*}=q(Z)/p^{*}(Z)$, we first claim that the RIPr satisfies $p^{*}(Z)\leq p_{\text{max}}(Z)$ up to $\mu$-nullsets. This is obvious if the RIPr belongs to ${\mathcal{P}}$, but in general it only belongs to ${\mathcal{P}}^{\circ\circ}$. In this case the claim follows by applying the following lemma with $f=p_{\text{max}}$.
@@
Let $f$ be a random variable such that $f\geq p$, $\mu$-almost surely, for all $p\in{\mathcal{P}}$. Then $f\geq p$, $\mu$-almost surely, for all $p\in{\mathcal{P}}^{\circ\circ}$.
@@
It is clear that $f\geq p$, $\mu$-almost surely, for all sub-probability densities $p$ dominated by some element of the convex hull of ${\mathcal{P}}$, and then also for all $p$ in the $\mu$-probability closure of this set. This was called ${\mathcal{P}}^{\prime\prime}$ in the proof of Theorem 4.7, where it was shown that ${\mathcal{P}}^{\prime\prime}$ is closed, convex, and solid. By Lemma 4.5, ${\mathcal{P}}^{\prime\prime}$ contains ${\mathcal{P}}^{\circ\circ}$, completing the proof. ∎
@@ 8 Summary
We established that under no assumptions on the composite null ${\mathcal{P}}$ and point alternative ${\mathsf{Q}}$, there exists a unique special e-variable $X^{*}$, the numeraire, which satisfies ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ for every other e-variable $X$, and is also log-optimal. Further, $X^{*}$ can be written as a likelihood ratio $d{\mathsf{Q}}/d{\mathsf{P}}^{*}$ for a unique (sub-)probability distribution ${\mathsf{P}}^{*}$, absolutely continuous with respect to ${\mathsf{Q}}$ and belonging to the bipolar ${\mathcal{P}}^{\circ\circ}$ of ${\mathcal{P}}$ (the effective null hypothesis). The numeraire is the only e-variable of this form. A strong duality theory establishes that ${\mathbb{E}}_{\mathsf{Q}}[\log X^{*}]$ equals the minimum relative entropy of ${\mathsf{Q}}$ to ${\mathcal{P}}^{\circ\circ}$, achieved by ${\mathsf{P}}^{*}$. This fully generalizes (to composite nulls) the fact that the likelihood ratio is the log-optimal e-variable for point nulls. We gave several sufficient conditions to identify and certify the numeraire, and showed that these were easy to verify in several nonparametric examples without a reference measure that were out of the reach of earlier methods. We also showed how to generalize our theory beyond log-optimality. Our work has implications for composite alternatives and for sequential testing using nonnegative supermartingales and ‘e-processes’ (sequential generalizations of e-variables). The latter topic introduces additional subtleties through dealing with filtrations and stopping times, a direction we are currently pursuing.

### v4

@@ 7 Universal inference is inadmissible
To compare the universal inference e-variable with the numeraire $X^{*}=q(Z)/p^{*}(Z)$, we first claim that the RIPr satisfies $p^{*}(Z)\leq p_{\text{max}}(Z)$ up to $\mu$-nullsets. This is obvious if the RIPr belongs to ${\mathcal{P}}$, but in general it only belongs to ${\mathcal{P}_{\textnormal{eff}}}$. In this case the claim follows by applying the following lemma with $f=p_{\text{max}}$.
@@
Let $f$ be a random variable such that $f\geq p$, $\mu$-almost surely, for all $p\in{\mathcal{P}}$. Then $f\geq p$, $\mu$-almost surely, for all $p\in{\mathcal{P}_{\textnormal{eff}}}$.
@@
It is clear that $f\geq p$, $\mu$-almost surely, for all sub-probability densities $p$ dominated by some element of the convex hull of ${\mathcal{P}}$, and then also for all $p$ in the $\mu$-probability closure of this set. This was called ${\mathcal{P}}^{\prime\prime}$ in the proof of Theorem 4.7, where it was shown that ${\mathcal{P}}^{\prime\prime}$ is closed, convex, and solid. By Lemma 4.5, ${\mathcal{P}}^{\prime\prime}$ contains ${\mathcal{P}_{\textnormal{eff}}}$, completing the proof. ∎
@@ 8 Summary
We established that under no assumptions on the composite null ${\mathcal{P}}$ and point alternative ${\mathsf{Q}}$, there exists a unique special e-variable $X^{*}$, the numeraire, which satisfies ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]\leq 1$ for every other e-variable $X$, and is also log-optimal. Further, $X^{*}$ can be written as a likelihood ratio $d{\mathsf{Q}}/d{\mathsf{P}}^{*}$ for a unique (sub-)probability distribution ${\mathsf{P}}^{*}$, absolutely continuous with respect to ${\mathsf{Q}}$ and belonging to the effective null hypothesis ${\mathcal{P}_{\textnormal{eff}}}$ associated with ${\mathcal{P}}$. The numeraire is the only e-variable of this form. A strong duality theory establishes that ${\mathbb{E}}_{\mathsf{Q}}[\log X^{*}]$ equals the minimum relative entropy of ${\mathsf{Q}}$ to ${\mathcal{P}_{\textnormal{eff}}}$, achieved by ${\mathsf{P}}^{*}$. This fully generalizes (to composite nulls) the fact that the likelihood ratio is the log-optimal e-variable for point nulls. We gave several sufficient conditions to identify and certify the numeraire, and showed that these were easy to verify in several nonparametric examples without a reference measure that were out of the reach of earlier methods. We also showed how to generalize our theory beyond log-optimality.
One natural direction for future work is to extend our theory to composite alternatives, using the method of mixtures or the plug-in method. A second direction is to extend our theory to the sequential setting using nonnegative supermartingales and e-processes (sequential generalizations of e-variables). See Ramdas et al. (2023) for an elaboration of the italicized terms and existing work. The second topic introduces additional subtleties through dealing with filtrations and stopping times, a direction we are currently pursuing.
Acknowledgments
We thank the referees for insightful comments and helpful suggestions.
@@ Appendix A A generalized Lebesgue decomposition
Appendix B Overview of finiteness properties
The table below gives an overview of the relation between absolute continuity properties and finiteness of the numeraire and relative entropies. Each cell shows the possibilities for the property or quantity indicated in the column header, under the assumption indicated for that row. In the table, $X^{*}$ is the numeraire, ${\mathsf{P}}^{*}$ is the RIPr, the relative entropy $H({\mathsf{Q}}\mid{\mathsf{P}}^{*})$ is in (3.1), and the maximal description gain $H({\mathsf{Q}}\mid{\mathsf{P}}^{*}\leadsto{\mathcal{P}_{\textnormal{eff}}})$ is defined in Remark 3.9. We specifically highlight the possibility that $H({\mathsf{Q}}\mid{\mathsf{P}})$ may be infinite for all ${\mathsf{P}}\in{\mathcal{P}}$, but finite or even zero for the RIPr ${\mathsf{P}}^{*}\in{\mathcal{P}_{\textnormal{eff}}}$; this happens, for instance, in Examples 4.4 and 4.9. Furthermore, although $H({\mathsf{Q}}\mid{\mathsf{P}}^{*}\leadsto{\mathcal{P}_{\textnormal{eff}}})=\infty$ if ${\mathsf{Q}}\not\ll{\mathcal{P}}$, the modified quantity in Theorem 3.13(iii) will be equal to zero for ${\mathsf{P}}={\mathsf{P}}^{*}$.
${\mathsf{Q}}(X^{*}<\infty)=1$ ${\mathsf{P}}^{*}\ll{\mathsf{Q}}$ ${\mathsf{Q}}\ll{\mathsf{P}}^{*}$ $H({\mathsf{Q}}\mid{\mathsf{P}}^{*})$ $H({\mathsf{Q}}\mid{\mathsf{P}}^{*}\leadsto{\mathcal{P}_{\textnormal{eff}}})$
${\mathsf{Q}}\not\ll{\mathcal{P}}$
no yes no =absent=$\infty$= ∞ =absent=$\infty$= ∞
${\mathsf{Q}}\ll{\mathcal{P}}$
[…]
${\mathsf{Q}}\ll{\mathcal{P}}$
$H({\mathsf{Q}}\mid{\mathsf{P}})<\infty$ for some ${\mathsf{P}}\in{\mathcal{P}}$
yes yes yes <absent<$\infty$< ∞ $=0$
${\mathsf{Q}}\ll{\mathcal{P}}$
$H({\mathsf{Q}}\mid{\mathsf{P}})=\infty$ for all ${\mathsf{P}}\in{\mathcal{P}}$
[…]
${\mathsf{Q}}\ll{\mathcal{P}}\ll\mu$ for some $\mu$
$H({\mathsf{Q}}\mid{\mathsf{P}})=\infty$ for all ${\mathsf{P}}\in{\mathcal{P}}$
${\mathcal{P}}$ is convex
[…]